OpenStudy (pottersheep):

Simplify i^17 + i^19 please

OpenStudy (pottersheep):

I got -i...but my answer sheet says -2i!

OpenStudy (anonymous):

i^1+i^3. i+-i is 0, so I'm not sure whether or not your answer sheet is right.

OpenStudy (pottersheep):

oo i got 0 too my bad

OpenStudy (pottersheep):

thanks!!

myininaya (myininaya):

\[i^{4 \cdot 4+1}+i^{4 \cdot 4+3}=i^1+i^3=i+(-i)=0\]